OpenStudy (anonymous):

Simplify by collecting the like Terms: 6 p - 3 p= 15 + 6

OpenStudy (anonymous):

The p terms can be added together: 6p - 3p = (6-3)p = 3p. Then add 15+6 like you normally would. Then you get 3p = 21.

OpenStudy (anonymous):

3p=21, p=7

OpenStudy (anonymous):

3p=21 p=7